OpenStudy (anonymous):

Thats not correct angela21793 3 7/8 + 5 5/24 = 9 1/12 according to the anwsers in the back of the book. how did they come up with that???

OpenStudy (anonymous):

You're\[\cos^{-1} \csc \cot \cot \tan \rho \theta \iota \pi \pi \leftarrow \int\limits_{\prod_{\prod_{\sum_{\sum_{\sum_{\eta \lambda \theta \ge \subseteq \theta =\supset}^{?}}^{?}}^{?}}^{?}}^{?}}^{?}\] incorrect.

OpenStudy (angela210793):

Juanita explained it i guess..i had forgotten to add 5 while multiplying lemme explain

OpenStudy (angela210793):